OpenStudy (anonymous):

Can someone help me do this equation... 3x+6y+4z=12 Section title: equation of a plane Chapter title: (lines and planes in space)

OpenStudy (anonymous):

i ll try

OpenStudy (anonymous):

okay thx in advance

OpenStudy (anonymous):

the standard form of the equation of a line in a plane is Ax+By=C

OpenStudy (anonymous):

z = -3/4 (x+2 y-4)

OpenStudy (anonymous):

how did you find the answer?

OpenStudy (anonymous):

Solving 3x + 6y + 4z = 12 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-6y' to each side of the equation. 3x + 6y + -6y + 4z = 12 + -6y Combine like terms: 6y + -6y = 0 3x + 0 + 4z = 12 + -6y 3x + 4z = 12 + -6y Add '-4z' to each side of the equation. 3x + 4z + -4z = 12 + -6y + -4z Combine like terms: 4z + -4z = 0 3x + 0 = 12 + -6y + -4z 3x = 12 + -6y + -4z Divide each side by '3'. x = 4 + -2y + -1.333333333z Simplifying x = 4 + -2y + -1.333333333z if u want z then z = -3/4 (x+2 y-4)

OpenStudy (anonymous):

thanks @gowtham89

OpenStudy (anonymous):

would you be able to do this...

OpenStudy (anonymous):

x=t+2 y=4t

OpenStudy (anonymous):

section: ^Lines and planes in space: a step by step procedure